OpenStudy (anonymous):

You are paid $8.25/hour. You work 35 hours/week for 3 weeks. Your involuntary deductions are FICA (7.65%), federal withholding (13%), and state withholding (9%). How much is your realized income? (1 point)$203.13 $866.25 $288.75 $609.39

OpenStudy (anonymous):

First figure out what your full salary is, before the taxes. To do this, multiply 35 and 3 to get 105 hours total. Multiply this number and the salary, 105*8.25=866.25. Now you have to add the taxes up to figure out how much is deducted. 7.65+13+9=29.65% Multiply that number and the salary to find out how much is deducted. A shortcut for this is to multiply the 1 minus the tax and the salary, to find the total salary, but either way will give the same answer. The tax written in decimal form is 0.2965. (866.25)(1-0.2965)=609.41